# SEED: Customize Large Language Models with Sample-Efficient Adaptation for Code Generation

### v2

@@ Abstract
Although Large Language Models (LLMs) have made significant progress in code generation, they still struggle with code generation tasks in specific scenarios. These scenarios usually necessitate the adaptation of LLMs to fulfill specific needs, but the limited training samples available in practice lead to poor code generation performance. Therefore, how to effectively adapt LLMs to new scenarios with few training samples is a major challenge for current code generation. In this paper, we propose a novel adaptation approach named SEED, which stands for Sample-Efficient adaptation with Error-Driven learning for code generation. SEED leverages the errors made by LLMs as learning opportunities, using error revision to overcome its own shortcomings, thus achieving efficient learning. Specifically, SEED involves identifying error code generated by LLMs, employing Self-Revise for code revision, optimizing the model with revised code, and iteratively adapting the process for continuous improvement. Experimental results show that, compared to other mainstream fine-tuning approaches, SEED achieves superior performance with few training samples, showing an average relative improvement of 54.7% in Pass@1 on multiple code generation benchmarks. We also validate the effectiveness of Self-Revise, which generates revised code that optimizes the model more efficiently compared to the code samples from datasets. Moreover, SEED consistently demonstrates strong performance across various LLMs, underscoring its generalizability.
Index Terms:
@@
Code Generation, Sample-Efficient Adaptation, Large Language Model.
I Introduction
@@
Code generation is an important technology that can improve the efficiency and quality of software development. Given the human requirement expressed in natural language, code generation allows machines to generate executable programs that satisfy this requirement. Code generation has been a research hot topic in the fields of artificial intelligence, software engineering, and natural language processing. Recently, code generation technologies have made significant advancements in both academia and industry [1, 2, 3]. In particular, large language models (LLMs) demonstrate great potential in code generation tasks [4, 5, 6, 7, 8, 9]. However, LLMs still face significant challenges in code generation for some specific scenarios or domains [10, 11].
For specific code generation scenarios, fine-tuning is an essential adaptation approach to ensure LLMs fulfill particular needs [12, 13, 14, 15]. However, in these specific scenarios, it is difficult to obtain sufficient training samples for fine-tuning LLMs, due to common reasons such as industry secrecy and scarcity of resources. For example, in safety-critical scenarios like aerospace, medical devices, transportation, and financial industries, the generated code must adhere to specific security specifications, and accessing relevant data is often extremely difficult due to high confidentiality and strict access control. Under the circumstance of limited data, mainstream fine-tuning approaches might not enable LLMs to achieve the desired code generation performance and may even lead to a degradation in model performance [16, 17, 18], as shown in Figure 1. Consequently, how to effectively adapt LLMs to specific scenarios with limited data available is a major challenge for code generation in practice.
@@
Figure 1: The performance of direct generation, fine-tuning, and our proposed SEED on MBPP dataset with few training samples. The numbers on the bars indicate the training samples used by different approaches.
@@
The mainstream fine-tuning approaches use a large number of samples gathered under specific scenarios for training [19]. They enable model to exhaustively learn the features present in these samples and thus adapt to the specific scenarios. However, they have two disadvantages. First, compelling LLMs to relearn the entire code data of these new scenarios is inefficient. Considering that LLMs are pre-trained on large-scale and diverse data, it’s reasonably assumed that they possess a certain level of general knowledge, lacking only particular information for application in specific scenarios. Second, when faced with insufficient data volume or data drift, model may learn certain undesirable features (such as inaccurate or irrelevant programming knowledge and patterns), thereby affecting its learning efficiency and negatively impacting its final performance.
@@
To overcome the disadvantages of these mainstream fine-tuning approaches, we take inspiration from the error-driven learning observed in humans. First, error-driven learning requires learners to identify their errors through testing. It helps learners to identify what they have mastered and what they still need to learn, allowing them to narrow the scope of learning and avoid wasting efforts on irrelevancies. Second, through error revision, learners can understand their deficiencies and make targeted improvements, thus enhancing learning efficiency and effectiveness. This motivates us to explore approaches to achieve sample-efficient adaptation of LLMs for code generation guided by error-driven learning.
@@
In this paper, we propose SEED, a Sample-Efficient adaptation based on Error-Driven learning for code generation. SEED aims to alleviate the problem of poor code generation performance of fine-tuning LLMs in scenarios with few training samples. Following the error-driven learning, our approach proceeds in four steps: ❶ Error Code Collection. We identify and collect error codes generated by LLMs, aiming to mine the weaknesses of LLMs. ❷ Automatic Code Revision. To obtain revisions of error codes in a low-cost way, we design Self-Revise to realize automatic revision leveraging information in the original dataset and code execution feedback. ❸ Model Optimization. We optimize the LLMs using the revised code, making them focus on learning the revision of these critical errors, thereby improving the learning efficiency of LLMs. ❹ Iterative Adaptation. We adopt an iterative strategy, which involves repeating the preceding three steps, to continuously optimize and improve the performance of LLMs. Extensive experimental results demonstrate the superiority and generalizability of SEED in the sample-efficient adaptation of LLMs for specific code generation scenarios.
@@
  • We introduce error-driven learning for LLMs. It utilizes revisions of model’s erroneous outputs for training and has higher training efficiency compared to using samples in the dataset.
@@
  • We propose a sample-efficient adaptation approach of LLMs for code generation, named SEED. Based on the principle of error-driven learning, SEED continuously optimizes model by identifying and revising error codes generated by LLMs, enabling effective adaptation to specific code generation scenarios with limited samples.
  • SEED outperforms other mainstream fine-tuning approaches on three public code generation benchmarks across various LLMs with few training samples.
II Motivation Example
Figure 2: A motivation example of SEED.
Figure 3: An overview of the proposed SEED and its differences from traditional fine-tuning approaches.
@@ II Motivation Example
Aligning LLMs with specific scenarios and addressing their unique challenges by learning samples in the dataset is difficult, especially when training samples are limited. We present a motivation example in Figure 2 to clarify the advantages of using error-driven learning in the LLMs adaptation process of code generation.
[…]
@@
Further, we explore the effectiveness of adopting error-driven learning from the perspective of model optimization. We consider the potential significant discrepancy between the model-generated output and the sample in the dataset. By learning the revisions of the model’s erroneous outputs, we can find more effective navigation in the optimization process. This might provide a shorter, smoother path to a good local minimum compared to learning from samples in the dataset, rather than attempting to direct it toward a distant area that may not align well with its existing knowledge or biases. We conduct the statistical analysis of the discrepancies in the model’s latent representations [^2] Specifically, on MBPP dataset, we obtain erroneous outputs of CodeGen-2B, revisions of the outputs, and samples in MBPP. We concatenate the requirements with their code, input them into CodeGen-2B, and extract the hidden representations from the model’s final layer. Then, we compute the Euclidean distances within the model’s representational space to quantify the disparities between these three elements.. The findings reveal that the average distance between the model’s erroneous outputs and the dataset’s samples is 12.35, whereas the average distance between the erroneous outputs and their revisions is significantly lower, at 6.39. These experimental results suggest that within the model’s representation space, revised codes are closer and similar to the erroneous output codes than the original code samples. This evidence lends support to our hypothesis of why the error-driven learning method is more efficient.
@@
Therefore, our work is determined to explore the use of error-driven learning to achieve a sample-efficient adaptation approach, aimed at enhancing the performance of LLMs in specific code generation scenarios.
III Methodology
@@
In this section, we present the detailed methodology of SEED. Specifically, given a code generation scenario/domain with a limited-sample training dataset $\mathcal{D}_{train}=\{(r,c)\}$, where each sample pair $(r,c)$ consists of an input requirement $r$ and an associated example of desired output code $c$. For a pre-trained LLM $\mathcal{M}_{\theta}$ with parameter $\theta$, we aim to adapt $\mathcal{M}_{\theta}$ to the specific scenario of $\mathcal{D}_{train}$. SEED achieves sample-efficient adaptation of LLMs through four steps: Error Code Collection (§III-A), Automatic Code Revision (§III-B), Model Optimization (§III-C), and Iterative Adaptation (§III-D). The overview of SEED and its differences from traditional fine-tuning are demonstrated in Figure 3.
III-A Error Code Collection
In this step, we systematically identify and collect erroneous output of LLMs using testing as criteria.
@@ III-A Error Code Collection
We employ rejection sampling [21] to draw error code samples from the distribution produced by $\mathcal{M}_{\theta}$. For each requirement $r\in\mathcal{D}_{train}$, we sample
$c^{\prime}\sim\mathcal{M}_{\theta}(r)~{}|~{}\neg f,$ (1)
@@
$\textsc{TestEval}(r,c^{\prime})=\left\{\begin{array}[]{ll}0,&\text{if }c^{\prime}\text{ fails }S_{r},\\ 1,&\text{otherwise},\end{array}\right.$ (2)
where $S_{r}$ is a suit of test cases under the requirement $r$ and is equipped by the code generation dataset. When collecting the error codes for test failures, we can keep the failed test cases and error messages simultaneously for further error diagnosis.
To gain insights into the propensity of $\mathcal{M}_{\theta}$ to make certain errors, it is advisable to select error code sample $c^{\prime}$ for which the model demonstrates relatively high confidence. Therefore, among multiple error codes collected for the same $r$, we select the one with the highest generation probability  [^3] We determine the probability of the generated code by averaging the probabilities of each generated token..
Figure 4: Illustration of automatic code revision
III-B Automatic Code Revision
In this step, we perform automatic revision for error codes using our Self-Revise approach. Based on the LLM $\mathcal{M}_{\theta}$ itself, Self-Revise revises the error code by providing the information in the original dataset and pointing out the error with code execution feedback. Our objective is to derive a revised code that is as close as possible to the error code. Note that although there is already a correct code $c$ in the dataset, it may differ significantly from the error code, as illustrated by examples (a), (b), and (c) in Figure 3. The pipeline of automatic code revision is shown in Figure 4.
Specifically, we leverage the following parts as the input of Self-Revise:
@@ III-B Automatic Code Revision
$T=\operatorname{Template}(r,g,c^{\prime},m,t)$ (3)
@@
Following previous work [22, 23], we use two settings for Self-Revise, i.e., fine-tuning (FT) and few-shot prompting (FSP), to get the model $\mathcal{M}_{\operatorname{Revise}}$ for revising error codes.
Self-Revise (FT) entails the process of fine-tuning $\mathcal{M}_{\theta}$ with a small number of samples for the purpose of automatic code revision. The training objective is to minimize the following loss function:
$L(\theta)=\sum_{i}l_{ce}(\mathcal{M}_{\theta}(T_{i}),c^{*}_{i})$ (4)
where $l_{ce}$ represents the standard cross-entropy loss, and we update the parameters initialized with $\mathcal{M}_{\theta}$ to obtain the model $\mathcal{M}_{\operatorname{Revise}}$ in Self-Revise (FT).
Self-Revise (FSP) adopts the few-shot prompting technique and leverages $k$ examples of $T_{i}$ and $c^{*}_{i}$ to construct the prompt $P$ for aligning $\mathcal{M}_{\theta}$ to automatic code revision. In Self-Revise (FSP), $\mathcal{M}_{\operatorname{Revise}}(\mathbf{\cdot})$ is defined as $\mathcal{M}_{\theta}(P\ ||\ \mathbf{\cdot}\ )$, where ||||| | denotes the concatenation operation.
@@
$c^{*}\sim\mathcal{M}_{\operatorname{Revise}}(T)~{}|~{}f.$ (5)
where $c^{*}$ is retained if $\textsc{TestEval}(r,c^{*})=1$ in Eq. (2), i.e., the revised code $c^{*}$ passes its test cases. We sample multiple times and it is sufficient if $\mathcal{M}_{\operatorname{Revise}}$ could correctly revise the error code once. To prevent $\mathcal{M}_{\operatorname{Revise}}$ from simply replicating the provided correct solution $g$, we exclude the output identical to $g$. Subsequently, for each requirement $r$, and select the version that is most similar to the error code among the remaining code revisions.
III-C Model Optimization

### v3

# Exploring Data-Efficient Adaptation of Large Language Models for Code Generation

@@ Abstract.
Although Large Language Models (LLMs) have made significant progress in code generation, they still struggle with code generation tasks in specific scenarios. These scenarios usually necessitate the adaptation of LLMs to fulfill specific needs, but the limited training data available in practice leads to poor code generation performance. Therefore, how to effectively adapt LLMs to new scenarios with few training data is a major challenge for current code generation. In this paper, we propose a novel adaptation approach named DEED, which stands for Data-Efficient adaptation with Error-Driven learning for code generation. DEED leverages the errors made by LLMs as learning opportunities, using error revision to overcome their own shortcomings, thus achieving efficient learning. Specifically, DEED involves identifying error code generated by LLMs, employing Self-Revise for code revision, optimizing the model with revised code, and iteratively adapting the process for continuous improvement. Experimental results show that, compared to other mainstream fine-tuning approaches, DEED achieves superior performance with few training data, showing an average relative improvement of 46.2% in Pass@1 on multiple code generation benchmarks. We also validate the effectiveness of Self-Revise, which generates revised code that optimizes the model more efficiently compared to the code samples from datasets. Moreover, DEED consistently demonstrates strong performance across various LLMs, underscoring its applicability.
@@
Code Generation, Data-Efficient Adaptation, Large Language Models
journal: TOSEMisbn: 978-1-4503-XXXX-X/18/06ccs: Software and its engineering Software creation and managementccs: Computing methodologies Artificial intelligence
1. Introduction
@@
Code generation is an important technology that can improve the efficiency and quality of software development. Given the human requirement expressed in natural language, code generation allows machines to generate executable programs that satisfy this requirement. Code generation has been a research hot topic in the fields of artificial intelligence, software engineering, and natural language processing. Recently, code generation technologies have made significant advancements in both academia and industry (Chen et al., 2021; Shen et al., 2022; Rozière et al., 2023; Dong et al., 2025a, 2023c). In particular, large language models (LLMs) demonstrate great potential in code generation tasks (Zheng et al., 2023; Nijkamp et al., 2023; Fried et al., 2022; Chen et al., 2023d; Zhang et al., 2023c; Jiang et al., 2023). However, LLMs still face significant challenges in code generation for adapting to specific scenarios or domains (Ahmed et al., 2024; Chen et al., 2023e).
@@
Figure 1. The performance of direct generation, fine-tuning, and our proposed DEED on MBPP dataset under the circumstance of limited data. The numbers on the bars indicate the training data used by different methods.
For specific code generation scenarios, fine-tuning is an essential adaptation method to ensure LLMs fulfill particular needs (Shi et al., 2023; Liu et al., 2024; Chakraborty et al., 2022; Ciniselli et al., 2022). However, in these specific scenarios, it is difficult to obtain sufficient training data for fine-tuning LLMs due to common reasons such as industry secrecy and scarcity of resources. For example, the programs in aerospace, medical devices, and transportation, where training samples are inherently scarce due to the scarcity of high-value industry data, high collection costs, and the typical targeting of long-tail demands. Under the circumstance of limited data, mainstream fine-tuning methods might not enable LLMs to achieve the desired code generation performance and may even lead to a degradation in model performance (Aghajanyan et al., 2021; Xu et al., 2021b; Zhang et al., 2022), as shown in Figure 1. In such data-scarce scenarios, only a small amount of data is available, but high code generation performance is still required, which is hard to solve but important in real-world applications. Consequently, how to effectively adapt LLMs to specific scenarios with limited data available is a major challenge for code generation in practice.
@@
The mainstream fine-tuning methods use a large number of samples gathered under specific scenarios for training (Xu et al., 2021a). They enable the model to exhaustively learn the features present in these samples and thus adapt to the specific scenarios. However, they have two disadvantages. First, compelling LLMs to relearn the entire code samples of new scenarios is inefficient. Considering that LLMs are pre-trained on large-scale and diverse samples, it’s reasonably assumed that they possess a certain level of general knowledge, lacking only particular information for application in specific scenarios. Second, when faced with insufficient data volume or data drift, the model may learn certain undesirable features (such as inaccurate or irrelevant programming knowledge and patterns), thereby affecting its learning efficiency and negatively impacting its final performance.
@@
To overcome the disadvantages of mainstream fine-tuning methods, we take inspiration from the error-driven learning observed in humans. 1) Error-driven learning requires learners to identify their errors through testing. It helps learners to identify what they have mastered and what they still need to learn, allowing them to narrow the scope of learning and avoid wasting efforts on irrelevancies. 2) Through error revision, learners can understand their deficiencies and make targeted improvements, thus enhancing learning efficiency and effectiveness. This motivates us to explore methods to achieve data-efficient adaptation of LLMs for code generation guided by error-driven learning.
@@
In this paper, we propose DEED, a Data-Efficient adaptation based on Error-Driven learning for code generation. DEED aims to alleviate the problem of poor code generation performance of fine-tuning LLMs in data-scarce scenarios. Following the error-driven learning, our method proceeds in four steps: ❶ Error Code Collection. We identify and collect error codes generated by LLMs, aiming to mine the weaknesses of LLMs. ❷ Automatic Code Revision. To obtain revisions of error codes in a low-cost way, we design Self-Revise to realize automatic revision leveraging information in the original dataset and code execution feedback. ❸ Model Optimization. We optimize the LLMs using the revised code, making them focus on learning the revision of these critical errors, thereby improving the learning efficiency of LLMs. ❹ Iterative Adaptation. We adopt an iterative strategy, which involves repeating the preceding three steps, to continuously optimize and improve the performance of LLMs.
We extensively evaluate our proposed DEED on five public code generation benchmarks using the test pass rate metric. Our results show that under the circumstance of limited data, DEED achieves significantly better performance across various LLMs compared to the mainstream adaptation approaches. Figure 1 illustrates part of this result. Specifically: 1) On five benchmarks, DEED achieves more than 27.1% relative improvement in the Pass@1 metric compared to the best-performing adaptation approaches. 2) The effectiveness of Self-Revise is validated. It produces revised code that better supports the optimization of the code generation model than the code samples provided in training dataset. 3) Our iterative adaptation process facilitates continuous enhancements in performance and maintains training stability. 4) DEED proves to be effective across various LLMs, demonstrating its broad applicability. These results highlight the potential of DEED in addressing the challenges of code generation in data-scarce scenarios.
@@
  • We propose that error-driven learning for LLMs adaptation is effective, i.e., utilizing revisions of LLMs’ erroneous outputs for training has higher learning efficiency than samples from the dataset.
@@
  • Based on the principle of error-driven learning, we propose a data-efficient adaptation method of LLMs for code generation, named DEED, which can effectively adapt the model to specific scenarios with few training data.
  • DEED outperforms the mainstream fine-tuning and prompting methods on five code generation datasets across various LLMs.
2. Preliminary Study
@@
Aligning LLMs with specific scenarios and addressing their unique challenges by learning samples in the dataset is difficult when training data are limited. We conduct a preliminary study to explore the effectiveness of using error-driven learning in the LLMs adaptation process of code generation.
@@
We consider the potential significant discrepancy between the model-generated output and the sample in the dataset. By learning the revisions of the model’s erroneous outputs, we can find more effective navigation in the optimization process. This might provide a shorter, smoother path to a good local minimum compared to learning from samples in the dataset, rather than attempting to direct it toward a distant area that may not align well with its existing knowledge or biases. We conduct a statistical analysis of the discrepancies in the model’s latent representations. Specifically, on MBPP (Austin et al., 2021) dataset, we obtain erroneous outputs of CodeGen-2B (Nijkamp et al., 2023), revisions of the outputs, and samples in MBPP. We concatenate the requirements with their code, input them into CodeGen-2B, and extract the hidden representations from the model’s final layer. Then, we compute the Euclidean distances within the model’s representational space to quantify the disparities between these three elements. The experimental results show that the average distance between the model’s erroneous outputs and the dataset’s samples is 12.35, whereas the average distance between the erroneous outputs and their revisions is significantly lower, at 6.39. These experimental results suggest that within the model’s representation space, revised codes are closer and similar to the erroneous output codes than the original code samples. This evidence supports our hypothesis of why the error-driven learning method is more efficient from the perspective of model optimization.
@@
Therefore, our work is determined to explore the use of error-driven learning to achieve a data-efficient adaptation method, aimed at enhancing the performance of LLMs in specific code generation scenarios.
Figure 2. An overview of the proposed DEED and its differences from traditional fine-tuning methods.
3. Methodology
@@
Given a code generation scenario/domain with a limited-sample training dataset $\mathcal{D}_{train}=\{(r,c)\}$, where each data pair $(r,c)$ consists of an input requirement $r$ and an associated example of desired output code $c$. For a pre-trained LLM $\mathcal{M}_{\theta}$ with parameter $\theta$, we aim to adapt $\mathcal{M}_{\theta}$ to the specific scenario of $\mathcal{D}_{train}$. Inspired by error-driven learning, we propose DEED to achieve data-efficient adaptation of LLMs. DEED consists of the following four steps: Error Code Collection (§3.1), Automatic Code Revision (§3.2), Model Optimization (§3.3), and Iterative Adaptation (§3.4). LLMs are well adapted after applying DEED, and their subsequent use (inference) remains unchanged compared to direct generation with LLMs, while improving the accuracy of one-time predictions without introducing any additional overhead. The overview of DEED and its differences from traditional fine-tuning are shown in Figure 2.
3.1. Error Code Collection
@@
In this step, we systematically identify and collect erroneous output of LLMs using testing as criteria. We employ rejection sampling (Casella et al., 2004) to draw error code samples from the distribution produced by $\mathcal{M}_{\theta}$. For each requirement $r\in\mathcal{D}_{train}$, we sample
(1) $c^{\prime}\sim\mathcal{M}_{\theta}(r)~|~\neg f,$
@@
(2) $\textsc{TestEval}(r,c^{\prime})\coloneqq\left\{\begin{array}[]{ll}0,&\text{if }c^{\prime}\text{ fails }S_{r},\\ 1,&\text{otherwise},\end{array}\right$.
where $S_{r}$ is a suit of test cases under the requirement $r$ and is equipped by code generation datasets. When collecting error codes for test failures, we can keep the failed test cases and error messages simultaneously for further error diagnosis.
To gain insights into the propensity of $\mathcal{M}_{\theta}$ to make certain errors, it is advisable to select error code sample $c^{\prime}$ for which the model demonstrates relatively high confidence. Therefore, among multiple error codes collected for the same $r$, we select the one with the highest generation probability, which is determined by the average log-probability of each token in the generated code.
(3) $\log P(c^{\prime}|r)=\frac{1}{|c^{\prime}|}\sum_{i}\log P(s_{i}|s_{<i},r),$
where $s_{i}$ is the $i$-th output token, $s_{<i}$ denotes the set of previous tokens, and $|c^{\prime}|$ indicates the length of the code.
3.2. Automatic Code Revision
In this step, we perform automatic revision for error codes using our Self-Revise method. Based on the LLM $\mathcal{M}_{\theta}$ itself, Self-Revise revises the error code by providing the information in the original dataset and pointing out the error with code execution feedback. Our objective is to derive a revised code that fixes the critical bug in the error code. As illustrated by examples (a), (b), and (c) in Figure 2, although there is already a correct code $(c)$ in the dataset, it may differ significantly from the error code $(a)$, leading to the critical bug being unclear. The pipeline of automatic code revision is shown in Figure 3. Specifically, we leverage the following parts as the input of Self-Revise:
@@ 3.2. Automatic Code Revision
(4) $T=\operatorname{Template}(r,g,c^{\prime},m,t)$
@@
Figure 3. Illustration of automatic code revision.
Following previous work (Zhang et al., 2023b; Dong et al., 2023b), we use two settings for Self-Revise, i.e., fine-tuning (FT) and few-shot prompting (FSP), to get $\mathcal{M}_{\operatorname{Revise}}$ for revising error codes.
Self-Revise (FT) entails the process of fine-tuning $\mathcal{M}_{\theta}$ with a small number of data for the purpose of automatic code revision. The training objective is to minimize $L(\theta)$:
(5) $L(\theta)=\sum_{i}l_{ce}(\mathcal{M}_{\theta}(T_{i}),c^{*}_{i})$
where $l_{ce}$ represents the standard cross-entropy loss, and we update the parameters initialized with $\mathcal{M}_{\theta}$ to obtain $\mathcal{M}_{\operatorname{Revise}}$ in Self-Revise (FT).
Self-Revise (FSP) adopts the few-shot prompting technique and leverages $k$ examples of $T_{i}$ and $c^{*}_{i}$ to construct the prompt $P$ for aligning $\mathcal{M}_{\theta}$ to automatic code revision. In Self-Revise (FSP), $\mathcal{M}_{\operatorname{Revise}}(\mathbf{\cdot})$ is defined as $\mathcal{M}_{\theta}(P\ ||\ \mathbf{\cdot}\ )$, where $||$ denotes the concatenation operation.
@@
(6) $c^{*}\sim\mathcal{M}_{\operatorname{Revise}}(T)~|~f.$
where $c^{*}$ is retained if $\textsc{TestEval}(r,c^{*})=1$ in Eq. (2), i.e., the revised code $c^{*}$ passes its test cases. We sample multiple times, and it is sufficient if $\mathcal{M}_{\operatorname{Revise}}$ could correctly revise the error code once. To prevent $\mathcal{M}_{\operatorname{Revise}}$ from simply replicating the provided correct solution $g$, we exclude the output identical to $g$. Subsequently, for each requirement $r$, select the version that is most similar to the error code among the remaining code revisions.
3.3. Model Optimization